OpenStudy (anonymous):

Solve for x. 5(2x - 1) = 6 x = 1/10 x = 11/10 x = 1/2

OpenStudy (anonymous):

11/10

OpenStudy (kropot72):

5(2x - 1) = 6 10x - 5 =6 10x = 11 x = 11/10